Given the two planes x - y + 2z = 0 and 3x + 2y - 6z + 4 = 0. Find a parametric equation for the intersection.


1
Expert's answer
2022-05-26T08:08:18-0400

If two planes intersect, they intersect in a line.

Setting "z=t," we can find the equation of the line.

"x-y+2z=0\\
\implies x-y+2t=0\\
\implies x-y=-2t\\"

"3x+2y-6z+4=0\\
\implies 3x+2y-6t+4=0\\
\implies 3x+2y=-4+6t"


"2x-2y=-4t\\
3x+2y=-4+6t\\
5x=-4+2t\\
x=-0.8+0.4t\\
y=x+2t=-0.8+0.4t+2t=-0.8+2.4t."


The parametric equation for the intersection:

"x=-0.8+0.4t,y=-0.8+2.4t,z=t."